OpenStudy (anonymous):

Solve 3/8(3+2x)+1≥13

OpenStudy (mathteacher1729):

Is that a 3/8 TIMES (other stuff) Or is it 3 OVER ( 8 times stuff , all under 3) ?

OpenStudy (anonymous):

is it: \[\frac{3}{8}(3+2x) + 1 \ge 13\] or \[\frac{3}{8(3+2x)} + 1 \ge 13\]

OpenStudy (anonymous):

it's 3/8 then (3+2x)+1≥13

OpenStudy (anonymous):

like your first example polpak

OpenStudy (anonymous):

Ok, so start by distributing the 3/8 to each term in the parens, then move your constant terms to the other side.

OpenStudy (anonymous):

ok...I multiply each number in parenthesis to 3 right?

OpenStudy (anonymous):

\[\frac{3}{8}(3 + 2x) = \frac{3}{8}(3) + \frac{3}{8}(2x)\]

OpenStudy (anonymous):

ok. so it would be 3/8(15x)+1≥13

OpenStudy (anonymous):

Nope.. What is \(\frac{3}{8}(3)\)

OpenStudy (anonymous):

i am really bad when fractions get involved .....9/24?

OpenStudy (anonymous):

\[\frac{3}{8} \times 3 = \frac{3}{8} \times \frac{3}{1} = \frac{3 \times 3}{8 \times 1} = \frac{9}{8}\]

OpenStudy (anonymous):

You need to practice your fractions. They will only become more important as time goes on.

OpenStudy (anonymous):

that's what I originally had but the way you put it was like I had to multiply 3 to both. so that makes sense. to do it that way.

OpenStudy (anonymous):

Ok, but then.. \[\frac{3}{8}(2x) = ?\]

OpenStudy (anonymous):

3/8(2x)= 3/8(2x/1) =6x/8

OpenStudy (anonymous):

or 6/8x

OpenStudy (anonymous):

Right. But now you cannot combine the 6x/8 and the 9/8. Since the 9/8 doesn't have an x.

OpenStudy (anonymous):

No, that's bad notation. It's either 6x/8 or it's \(\frac{6}{8}x\) or (6/8)x It is not 6/8x

OpenStudy (anonymous):

so now I have to get x alone right?

OpenStudy (anonymous):

Right. So you have: \[\frac{9}{8} + \frac{6}{8}x + 1 \ge 13\]

OpenStudy (anonymous):

that's what I was trying to text but I don't know how to make a fraction the right way on here.

OpenStudy (anonymous):

and you want to move all the non-variable terms to one side and the variable ones on the other.

OpenStudy (anonymous):

so I multiply by 6/8 right ?

OpenStudy (anonymous):

or 8/6?

OpenStudy (anonymous):

No. You subtract off those non-variable terms.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Subtract 1 from both sides. Subtract 9/8 from both sides.

OpenStudy (anonymous):

If you want to rid yourself of the fraction you can multiply both sides by 8 if you like, though if you were gonna do that it would have been easier from the beginning.

OpenStudy (saifoo.khan):

hey rachel, im back!

OpenStudy (anonymous):

Yay Saifoo!!!!

OpenStudy (saifoo.khan):

\[\color{red}{\text{so how should v start??}}\]

OpenStudy (saifoo.khan):

:D

OpenStudy (anonymous):

check your messages

OpenStudy (saifoo.khan):

yay, im there